OpenStudy (anonymous):

Will Fan and Medal, Given the functions f(n) = 25 and g(n) = 3(n − 1), combine them to create an arithmetic sequence, an, and solve for the 12th term.

OpenStudy (anonymous):

Given the functions f(n) = 25 and g(n) = 3(n − 1), combine them to create an arithmetic sequence, an, and solve for the 12th term. an = 25 − 3(n − 1); a12 = −11 an = 25 − 3(n − 1); a12 = −8 an = 25 + 3(n − 1); a12 = 58 an = 25 + 3(n − 1); a12 = 61

OpenStudy (anonymous):

I dont understand how to combine them exactly.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@k_lynn